if a+b+c= 3 and ab=2 then find value of (i)a^2+b^2 (ii) a-b
 

Grade:9

1 Answers

ankush rathee
6 years ago
(i) a^2+b^2=(a+b)^2-2ab =(3-c)^2-4 =9+c^2-6c -4 =c^2-6c+5 = (c-1)(c-5)(ii) let (a-b) be X So (a-b)^2 =a^2+b^2-2ab=c^2-6c+5-4=c^2-6c+1So a-b= root (c^2-6c+1)